01:30 some thin clouds on the horizon. night might not be photometric.

02:18 attempted to move telescope to zenith, but encountered HBS problem.

02:28 moving to Skyflat position (~5-15 deg off of zenith).

02:30 still see bands of thin clouds, low on the S horizon. night is likely not quite photometric.

02:31 turning on LBC...test bias frame shows some banding on the red side.

02:40 taking U/Y test skyflat (scale=1): saturated on blue side.

02:44 taking U/Y test skyflat (scale=0.5): 24k on blue and 44k on red.

02:45 taking U/Y skyflats (scale=0.7, 5 frames): first frame has 24k on blue and 35k on red. fifth frame has 14k blue and 16k on red.

02:51 taking Us/Y skyflats (scale=2.0, 5 frames): first frame has 24k on blue and 19k on red. 13k in last blue frame.

02:56 taking B/r skyflats (scale=2.5, 5 frames): 28.5k on blue side and 34k on red side in first frame. 16k in last blue image.

03:03 moving to standard star field...and then waiting for sky to darken...

03:15 focusing....

03:27 running std star OB: BD+26_UY. (uses Us)

03:37 running a modified std star OB: longer exposure time and U filter in this OB.

03:42 moving to SA104...

03:44 focusing on std field...

03:49 running std star OB (SA104_UsY)...

03:54 running std star OB (SA104_UY)...

04:01 moving to EGS_Y_1b and focusing...

04:07 running EGS_Y_1b OB: seeing of ~3.2pix and ~5.2pix in red and blue, respectively (in first image).

04:25 rotator fault on LBCB during 4th exposure of OB. turning LBC off and on to solve issue. this did not work, so now running lbckill/lbcstart. LBC is now starting up...

04:33 moving to EGS_Y_1a and focusing...

04:38 starting EGS_Y_1a OB.

05:03 refocusing...

05:06 running EGS_Y_1b OB.

05:32 refocusing...

05:37 starting EGS_Y_1b_last3 OB (just last 3 dither positions of EGS_Y_1b).

05:52 moving to Bootes b36r...telescope or enclosure fault! (unknown telescope-related fault)

06:09 finally at Bootes....now focusing...

06:13 starting priority5_PA0: b36r: seeing of ~0.6" in the red and ~0.85" in the blue (in first image). guiding jump in image lbcr.20080603.062434.fits. seeing of ~3pix in lbcr.20080603.062911.fits. John Hill called in to say that the focus appeared to be bad on the red guider. Our images looked good, so we did not apply an auto-focus adjustment, as he recommended.

06:39 moving b56r....and then refocusing....

06:45 running priority4_PA180: b56r: small guiding jump in first image (seeing is ~3pix in red and ~5pix in blue).

07:11 refocusing...

07:13 running on priority1_PA180: b24r: seeing is 4.7pix and 3.1pix in frame lbcb.20080603.071909.fits and lbcr.20080603.071905.fits.

07:38 refocusing...

07:44 running priority5_PA0: b25r: seeing of 3.3-3.4 pix in red (lbcr.20080603.075407.fits) and >5pix in blue (lbcb.20080603.075410.fits).

08:08 moving to SA107...refocusing...

08:08 running SA107_UsY OB...seeing of ~0.75 in Y-band.

08:19 running SA107_UY OB...problem with red channel communication (we're at 45deg elevation)....turning LBC off and on, which did not work.

08:34 OB constructed for SA110, which is at higher elevation. moving SA110...then we will attempt to bring camera back online.

08:36 turning camera off and on again (we're now at elevation of 56deg)...LBC came back online!

08:41 focusing...

08:44 starting SA110_UY OB...

08:49 running SA110_Br...

08:54 moving to Herc_4 and focusing...

08:59 observing Herc_4 in B-only mode (only 2 frames)...seeing is ~5.5pix in B.

09:11 moving to Herc_5 and focusing...

09:16 observing Herc_5 in B-only mode (only 3 frames)...seeing is ~6pix in B.

09:33 moving to Herc_9 and focusing...

09:36 observing Herc_9 in B-only mode (only 3 frames)...seeing is >6.5pix in B.

09:54 moving to Herc_10 and focusing...

09:59 starting Herc_10 in B-only mode (only 4 frames)...seeing is ~6.2pix in B.

10:22 moving to Herc_8 and refocusing...

10:27 starting Herc_8 in B-only mode (only 3 frames)....

10:45 moving to SA110 and refocusing...

10:57 starting SA110_UsY...

11:01 starting SA110_UY...

11:07 starting SA110_Br...

11:11 moving to LDS749B spectro standard and focusing...

11:16 starting LDS749B_UsY...

11:21 starting LDS749B_UY...

11:26 moving to skyflat position...

11:28 test skyflat in B/r (scale=5)

11:30 test skyflat in B/r (scale=5)

11:32 skyflat OB in B/r (scale=5)...5 frames.

11:38 test skyflat in Us/Y (scale=2)

11:40 skyflat OB in Us/Y (scale=3)...5 frames.

11:44 test skyflat in U/Y (scale=1)

11:45 skyflat OB in U/Y (scale=1)...5 frames.

11:50 closing dome...then moving telescope back to zenith, in order to take bias frames.

12:00 running 25Bias_Bino OB...OB stopped due to warning from LBC-red. connection problems again. turning LBC off and on (at zenith) recovers...

12:05 running 25Bias_Bino OB...

At sunrise, looks clear outside. Hopefully the standards will be useful. Fairly productive night although seeing degenerated later.

-- XiaohuiFan - 03 Jun 2008
Topic revision: r3 - 03 Jun 2008, XiaohuiFan
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback